OpenStudy (anonymous):

Other method?!? Someone asked me, I said there is not! What do you think? \[\lim_{x \rightarrow 0} \frac{e-(1+x)^{\frac{1}{x}}}{x}\] No Hopital, no series expansion

OpenStudy (anonymous):

not sure how you would expand in any case

OpenStudy (anonymous):

@satellite73

OpenStudy (ikram002p):

well convert to x goes to infinity as of 1/x

OpenStudy (anonymous):

@ikram002p and then?

OpenStudy (ikram002p):

\(\large \lim_{x \rightarrow \infty } \frac{e-(1+\frac{1}{x} )^{x}}{\frac{1}{x}}\) haha seems also LH :P

OpenStudy (anonymous):

LH can't it is infinity over zero

OpenStudy (anonymous):

11 hours ago ??!!! I posted it right now :))

OpenStudy (anonymous):

@Catch.me it's 0/0

OpenStudy (anonymous):

how did you get e-(1+x)^1/x = 0 ??

OpenStudy (ikram002p):

mmm so \(\large \lim_{x \rightarrow \infty } x (e-(1+\frac{1}{x} )^{x} )\) ?

OpenStudy (anonymous):

@Catch.me It is well known\[\lim_{x \rightarrow 0} (1+x)^{\frac{1}{x}}=\lim_{x \rightarrow \infty} (1+\frac{1}{x})^{x}=e\]

OpenStudy (ikram002p):

hehe so nw we have (infinity . 0) mmm can we think of squeeze theorem?

OpenStudy (anonymous):

That may work also, I don't know :-) Give it a try

OpenStudy (ikram002p):

im squeezing my head o.o

OpenStudy (anonymous):

I got it thanks @mukushla

OpenStudy (anonymous):

yes here: http://www.physicsforums.com/showthread.php?t=250172

OpenStudy (anonymous):

and so many other proofs, just google it ;)

OpenStudy (anonymous):

i did that

OpenStudy (anonymous):

OK then it is 0/0 why don't you use the quotient rule.

OpenStudy (anonymous):

quotient rule derivative? well it's Hopital, I look for another one

OpenStudy (anonymous):

I'll go back to this later :-) thank you all

ganeshie8 (ganeshie8):

*